OpenStudy (anonymous):

What is the equation of this problem? Victoria has twice as much money as Jacob. Jacob has $10 less than Jesse. Together they have $78. How much money does each have?

OpenStudy (anonymous):

Let jesse be x

OpenStudy (anonymous):

How much money does each have

OpenStudy (tkhunny):

Setting up the equation or equations is the part of the question where you will actually learn something. It is of little additional consequence if you can solve it. Setting it up remains for you to accomplish. Rule #1 - Name Stuff. Question #1 - Name what? Answer #1 - What does it want? Name that. "How much money does each have?" V = amount of money had by Victoria J = amount of money had by Jacob S = amount of money had by Jesse (J was already taken) Awesome. You are now ready to start translating the problem statement information into equations! Go!!

OpenStudy (anonymous):

Let jacob be (x-10)

OpenStudy (anonymous):

Let victoria be 2(x-10)

OpenStudy (anonymous):

Sum them all together. 2(x-10) +x-10+x=78

OpenStudy (anonymous):

done?

OpenStudy (tkhunny):

"Victoria has twice as much money as Jacob." V = 2J "Jacob has $10 less than Jesse." J = S - 10 "Together they have $78." V + J + S = 78 In this way, we have translated DIRECTLY from the words of the problem statement to useful equations. It was not required to do half the problem in your head. It's all written down before your eyes. The rest will test your algebra skills.

OpenStudy (anonymous):

^ right